# Distributed acoustic sensing for ocean applications

### v2

@@ Abstract
Extensive monitoring of acoustic activities is important for many fields, including biology, security, oceanography, and Earth science. Distributed acoustic sensing (DAS) is an evolving technique for continuous, wide-coverage measurements of mechanical vibrations, which is suited to ocean applications. DAS illuminates an optical fiber with laser pulses and measures the backscattered wave due to small random variations in the refractive index of the material. Specifically, DAS uses coherent optical interferometry to measure the phase difference of the backscattered wave from adjacent locations along the fiber. External stimuli, such as mechanical strain due to acoustic wavefields impinging on the fiber-optic cable, modulate the backscattered wave. Hence, the differential phase measurements of the optical backscatter are proportional to the underlying physical quantities of the surrounding wavefield. Continuous measurement of the backscattered electromagnetic signal provides a distributed sensing modality that extends spatially along the fiber. We provide an overview of DAS technology and detail the underlying physics, from electromagnetic to mechanical and eventually acoustic quantities. We explain the effect of DAS acquisition parameters in signal processing and show the potential of DAS for sound source detection on data collected from the Ocean Observatories Initiative Regional Cabled Array https://doi.org/10.58046/5J60-FJ89.
@@ I Introduction
Light propagating in an optical fiber undergoes backscattering due to small random variations in the refractive index of the material. Fiber-optic sensing technology is based on the phase modulation of the backscattered light traveling in an optical fiber due to external stimuli such as temperature gradients or mechanical strain. Distributed sensing transforms a fiber cable into a very long, dense sensor arraygrattan2000fiber that can extend over hundredswaagaard2021realtime or even thousands of kilometers,mazur2024real ; ronnekleiv2025range offering unprecedented sensing coverage with rapidly advancing technology. Moreover, fiber-optic sensing makes temperature and vibration measurements in harsh environments accessible due to its robustness to electromagnetic interference, high pressure, and temperature.Miah2017 ; Lindsey2021
@@
Distributed acoustic sensing (DAS) is a fiber-optic sensing technology, which illuminates an optical fiber with laser pulses and measures the phase of the backscattered wave along the fiber.Hartog2017 A fiber-optic cable that is in mechanical contact with the surrounding medium will be stressed by external mechanical vibrations. Strain, i.e., the relative deformation of the fiber under stress, alters the distance traveled by the backscattered light changing its phase.KennettBook DAS records the phase of the backscattered light to estimate the strain (or strain rate) along the fiber and infer the displacement (or velocity) of the mechanical vibration.masoudi2016contributed ; yang2022sub Elongated optical fibers are mostly sensitive to the axial component of the induced strain.martin2021introduction ; williams2024toward Consequently, without a tangled cable layout,kuvshinov2016interaction ; ning2018multicomponent DAS exhibits directional sensitivity to the measured mechanical wavefield.fang2023directional Given that stress and strain are related through the elastic modulus of the solid material according to Hooke’s law,Brekhovskikh1994 both the fiberglass coatingbudiansky1979pressure and the deployment conditions of the cable, e.g., burial depth and seafloor type,sladen2019distributed affect the strain sensitivity of the fiber optic cable to the applied stress.
@@
DAS has gained momentum in geophysics as a cost-effective, high-resolution method for seismic monitoring and seismic tomography.mateeva2014distributed ; zhan2020distributed ; Lindsey2021 Applications range from near-surface seismic monitoring of ambient noisedou2017distributed ; saengduean2024multi to borehole measurements of hydraulic fracturingjin2017hydraulic ; chien2023automatic and monitoring the sea state dynamics with submarine cables.lindsey2019illuminating ; lin2024measurement ; chien2025calibration Recently, the potential of DAS technology has been leveraged for monitoring the underwater acoustic wavefield with promising results in the identification and localization of low-frequency sound sources such as blue and fin whale vocalizationsbouffaut2022eavesdropping ; landro2022sensing ; wilcock2023distributed and ship noise.rivet2021preliminary ; shao2025tracking ; paap2025leveraging Monitoring the ocean soundscape at low frequencieswilcock2014sounds is key to understanding the impact of anthropogenic noiseandrew2011long ; chapman2011low ; mcdonald2006increases and climate changegavrilov2008long on ecology,gavrilov2012steady security, and marine exploration.miksis2016low So far, the characterization of the ocean soundscape has been based on recordings from hydrophones and hydrophone arrays with limited spatial coverage.mcdonald2006increases ; chapman2011low DAS converts the length of the fiber cable to an extensive linear aperture, which is especially useful for localization of low-frequency sound sources (below 50-100 Hz).
@@
Despite the increasing interest in DAS technology for high-resolution observations of acoustic signals, the current applications in ocean acoustics are limited to the aforementioned low-frequency wavefields, i.e., whale calls, ship noise, and ocean swell. The absence of ocean acoustic observations at higher frequencies is related to the acquisition parameters of the current DAS systems, which are set with the objective of a sufficient signal-to-noise ratio. Nevertheless, calibration experiments with controlled transmitted signalsdouglass2023distributed ; shen2024high and at-sea experiments for underwater communicationspotter2024first indicate that DAS detection capability is not limited to low-frequency sounds. This paper links the acquisition parameters with signal processing and provides insight for improved observations of acoustic signals.
@@
After trend removal and band-pass filtering to the frequency range of interest, the data are commonly analyzed with interferometric methods,saengduean2024multi $f$-$k$ decomposition, and delay-and-sum beamforming.nasholm2022array Given the high dimensionality of DAS data, machine learning provides compelling methods for dimensionality reduction and feature clustering.chien2023automatic
@@
In this paper, we provide an overview of DAS technology. We detail the mathematical derivation of the underlying physics, from electromagnetic to mechanical and eventually acoustic quantities in Section II. The mathematical analysis highlights the physical principles of DAS in a generic manner, aiming to inform the choice of data acquisition parameters rather than to detail the technological specificity of different systems, which are evolving rapidly. Section III conforms the spatial sensitivity of DAS to that of an array of spatially distributed sensors of a sizeable aperture. Section IV analyzes publicly available data for sound source detection based on the phase of the acoustic particle displacement. Finally, Section VII concludes the paper.
@@
$z$ axial distance $t$ time $\sin\theta$ directional of impinging wave $\widetilde{[\cdot]}$ complex quantity
$L_{f}$ interrogated fiber length $\tau_{\text{max}}$ two-way travel time along fiber length $\theta$ polar angle $[\cdot]^{*}$ complex conjugation
$L_{p}$ light pulse length $T_{p}$ light pulse duration (width) $k_{z}$ wavenumber of acoustic wave along the axial dimension $\widehat{[\cdot]}$ estimated quantity
$L_{g}$ gauge length $\Delta\Phi$ differential phase of backscattered light $[\cdot]_{z}$ projection to the fiber’s axis
$L_{h}$ channel distance - distance between spatial samples $\tau_{\text{max}}$ distance between temporal samples $[\cdot]_{a}$ quantity related to the acoustic field
$f_{s}$ sampling frequency (spatial sampling) $f_{r}$ pulse repetition frequency (temporal sampling) [][]delimited-[]delimited-[][$\cdot$]\!*\![$\cdot$][ ⋅ ] ∗ [ ⋅ ] convolution
Figure 1: (color online) Schematic representation of scattering mechanisms due to a light pulse of frequency $f_{p}$ and wavelength $\lambda_{p}$ propagating through an optical fiber. In single-mode fibers, $f_{p}$ is typically about 200 THz ($\lambda_{p}\approx$ 1550 nm). External stimuli, such as an increase in temperature $T$ or strain $\epsilon$, affect the respective scattering response as indicated by the arrows. The horizontal axis represents the wavelength $\lambda$ or, in reverse direction the frequency $f$ of the backscattered light. Stokes processes refer to an up-shift of the wavelength of the scattered light $\lambda>\lambda_{p}$, whereas anti-Stokes to a down-shift $\lambda<\lambda_{p}$. The Stokes radiation for Raman scattering is relatively insensitive to temperature and has a lower intensity than the corresponding anti-Stokes radiation due to lower scattering efficiency at lower frequencies.Ukil2011
@@ II Distributed Acoustic Sensing Technology
DAS measures the effect of external mechanical vibrations on the backscattering of laser light traveling along an optical fiber. This section details the DAS technology by (II.1) providing an overview of the light scattering mechanisms within a fiber optic cable, (II.2-II.3) deriving the mathematical expression for the backscattered wave under monochromatic excitation, (II.4) explaining the detection method of the backscattered light, (II.5) showing how the measurand is related to strain from external mechanical stress, and finally (II.6) explaining the typical acquisition parameters and the data structure. Table 1 provides a list of the important parameters for the analysis.
@@ II.1 Scattering mechanisms
An optical fiber is a thin thread of silica glass with a core of dielectric material, clad with a lower refractive index material. The difference in refractive index between the core and the cladding creates an optical waveguide for light transmitted in the fiber. Optical fibers are designed to have a core with a constant refractive index such that light transmitted into the fiber travels undisturbed along the fiber. However, in practice, small density variations in the core material due to impurities or the manufacturing procedure introduce small random variations in the refractive index along the optical fiber, which cause scattering of the propagating electromagnetic wave.KennettBook Hence, the refractive index along the fiber’s axis can be described as
$n(z)=n_{f}+\Delta n(z),$ (1)
where $n_{f}$ is the nominal refractive index of the optical fiber’s core material and $\Delta n(z)$ describes the localized fluctuation of the refractive index. The refractive index determines the speed of light in the optical fiber as $c_{n}\approx c/n_{f}$, where $c\approx 3\times 10^{8}$ m/s is the speed of light in vacuum.
The main scattering mechanisms are Rayleigh, Brillouin, and Raman scattering.masoudi2016contributed Rayleigh scattering is an elastic process caused by local scatterers with dimensions much smaller than the wavelength that does not involve energy transfer to the scatterer and does not shift the frequency of the incident wave. Brillouin and Raman scattering are inelastic processes caused by changes in the vibrational states in a lattice and molecular level, respectively, that involve energy transfer to the scatterer and, as a result, shift the frequency of the incident wave.Cotter1983 ; Shiota1992 Brillouin scattering can be leveraged for both strain and temperature measurements,Bao2011 whereas Raman scattering can be leveraged for distributed temperature sensing.Ukil2011 Rayleigh scattering exhibits the highest intensity and is principally exploited for DAS measurements.KennettBook Figure 1 depicts the effect of external stimuli such as temperature and strain to elastic and inelastic optical scattering mechanisms [Grattan2000, , Ch. 3.8].
Figure 2: (color online) Schematic representation of active optic sensing for passive monitoring of external vibrational fields. (a) A single-mode fiber of total length $L_{f}$ and a small cross-section constructed with randomly distributed impurities in the core material. (b) The fiber is interrogated with a laser pulse $E_{p}$ of duration $T_{p}$. (c) As the probe pulse travels along the fiber at light speed $c_{n}$, a point scatterer located at $z_{s}$ causes part of the incident energy $E_{i}$ to propagate back to the interrogator. (d) The backscattered wave $E_{s}$ from a single point scatterer is received around a time instant corresponding to the two-way time-of-flight between the laser source and the scatterer with a duration equal to the probe pulse. (e) An external vibration applies stress to the fiber, which deforms its shape. Under strain, the scatterers’ positions are slightly perturbed. (f) A scatterer’s displacement introduces a proportional delay to the backscattered wave. (g) The total backscattered field results from the superposition of the backscattered light from all illuminated scatterers along the fiber. (h) The time-of-flight $\tau$ of the received total backscattered signal $E_{b}$ is linearly related to the axial distance along the fiber. Cross-correlating the total backscattered signal received from adjacent fiber locations separated by the gauge length $L_{g}$, provides an estimate of the axial strain along the fiber due to external vibrations; see Sec II.5.
@@ II.2 Rayleigh backscattered wave
DAS technology employs an optoelectronic system, referred to as the interrogator unit, to transmit a light pulse into the optical fiber and detect the power of the backscattered electric field due to the interaction of the transmitted pulse with the impurities along the fiber.Hartog2017
The most common single-mode optical fibers are designed with a cross-section of micro-metric dimensions such that only a single mode (transverse mode) of electromagnetic wave propagation is supported without modal dispersion. For a typical single-mode optical fiber, a minimum attenuation of $0.18$ dB/km is achieved for a light wavelength of $1550$ nm [Hartog2017, , Fig. 2.3], i.e., at a frequency of $200$ THz. Hence, a narrowband laser pulse at this frequency would sustain a minimum transmission loss.Miah2017 ; Hartog2017
Multimode optical fibers have a core with a larger cross-section (at least an order of magnitude larger than single-mode fibers), which guides multiple electromagnetic modes simultaneously [Hartog2017, , Sec. 2.1]. Modal dispersion of the transmitted light as it propagates along the fiber eliminates its phase coherence and polarization information. Consequently, multimode optical fibers do not support highly sensitive coherent detection methods resulting in reduced detection range and resolution. However, light emitting diodes (LED) can be used for light transmission in multimode fibers instead of highly coherent laser sources required in single-mode fibers, resulting in cost-effective systems [Grattan2000, , Ch. 1]. Despite the increased cost, coherent single-mode fiber systems are the preferred solution for DAS due to the higher sensitivity.
Figure 2 depicts the mechanism of active optic sensing in single-mode fibers and indicates its utilization for passive sensing of external vibrations. In mathematical terms, the probe pulse [Fig. 2(b)] of duration $T_{p}$, at radial frequency $\omega$ and initial phase $\phi_{0}$, generated by the laser source at $z=0$ is
$E_{p}(t)=p\left(\frac{t}{T_{p}}\right)\cos\left(\omega t+\phi_{0}\right),$ (2)
where the temporal window $p$ is defined by the apodization function (wavelet) $w$ as
@@
Consider the impulse response received at a reference location $z_{0}=0$ as a function of the travel time $\tau$ due to an ideal point scatterer at $z_{s}$ as the two-way Green’s function
@@
where $\delta$ denotes the Dirac delta function, * is the convolution operator and $g_{i}$ and $g_{b}$ denote the Green’s function for the incident and the backscattered wave at location $z_{s}$ reached at travel time $t_{s}$ due to an impulse source at $z_{0}=0$ and $t_{0}=0$, respectively,
$g_{i}(t_{s},z_{s}|t_{0},z_{0})\!=\!\delta\!\left(\!t_{s}\!-\!t_{0}\!-\frac{z_{s}\!-z_{0}}{c_{n}}\!\right)\!=\!\delta\!\left(t_{s}-\frac{z_{s}}{c_{n}}\right),$ (5)
$g_{b}(t_{0},z_{0}|t_{s},z_{s})\!=\!\delta\!\left(\!t_{0}-\!t_{s}\!-\!\frac{z_{0}-z_{s}}{c_{n}}\!\right)\!=\!\delta\!\left(\frac{z_{s}}{c_{n}}-t_{s}\right).$
More generally, the backscattered wave [Fig. 2(d)] from a point scatterer at $z_{s}$ with scattering strength $s(z_{s})$ due to the excitation pulse (2) is
@@
External vibrations, such as an impinging acoustic wavefield at angle of incidence $\theta$ characterized by a particle displacement $\mathbf{u}$, apply stress to the fiber, which deforms its shape [Fig. 2(e)]. Under strain, any scatterer’s position is slightly perturbed by $\Delta z$, resulting in a proportional delay at the associated backscattered signal, while the scattering strength $s(z_{s})$ remains the same [Fig. 2(f)]:

### v3

# Overview of distributed acoustic sensing: Theory and ocean applications

@@ Abstract
Extensive monitoring of acoustic activities is important for many fields, including biology, security, oceanography, and Earth science. Distributed acoustic sensing (DAS) is an evolving technique for continuous, wide-coverage measurements of mechanical vibrations, which is suited to ocean applications. DAS illuminates an optical fiber with laser pulses and measures the backscattered wave due to small random variations in the refractive index of the material. External stimuli, such as mechanical strain due to acoustic wavefields impinging on the fiber-optic cable, modulate the backscattered wave. Continuous measurement of the backscattered signal provides a distributed sensing modality of the impinging wavefield. Considering the potential use of existing telecommunication fiber-optic cables deployed across the oceans, DAS has emerged as a promising technology for monitoring the underwater soundscape. This review presents advances in DAS in the last decade and details the underlying physics from electromagnetic to mechanical and eventually acoustic quantities. To guide the use of DAS for ocean applications, the effect of DAS acquisition parameters in signal processing is explained. Finally, DAS is demonstrated on data from the Ocean Observatories Initiative Regional Cabled Array for the detection of sound sources, such as whales, ships, and earthquakes.
@@ I Introduction
Light propagating in an optical fiber undergoes backscattering due to small random variations in the refractive index of the material. Fiber-optic sensing technology is based on the phase modulation of the backscattered light traveling in an optical fiber due to external stimuli such as mechanical strain or temperature changes. Distributed sensing transforms a fiber cable into a very long, dense sensor arraygrattan2000fiber that can extend over hundredswaagaard2021realtime or even thousands of kilometers,mazur2024real ; ronnekleiv2025range offering unprecedented sensing coverage with rapidly advancing technology. Moreover, fiber-optic sensing makes temperature and vibration measurements in harsh environments accessible due to its robustness to electromagnetic interference, high pressure, acidity, and temperature.Miah2017 ; Lindsey2021
@@
Distributed acoustic sensing (DAS) is a fiber-optic sensing technology that illuminates an optical fiber with laser pulses and measures phase differences of the backscattered wave along the fiber.Hartog2017 External mechanical vibrations will stress a fiber-optic cable that is in mechanical contact with the surrounding medium. Strain, i.e., the deformation of the fiber under stress, alters the two-way distance traveled by the laser pulse, changing the phase of the backscattered light.KennettBook DAS records the differential phase of the backscattered light to estimate the strain (or strain rate) along the fiber, which is in turn related to the displacement (or velocity) of the mechanical vibration.masoudi2016contributed ; yang2022sub Elongated optical fibers are mostly sensitive to the axial component of the induced strain.martin2021introduction ; williams2024toward To overcome the effect of the directional sensitivity of DAS to the measured mechanical wavefieldfang2023directional , helically wound fiber-optic cables have been proposed.kuvshinov2016interaction ; ning2018multicomponent Given that stress and strain are related through the elastic modulus of the solid material according to Hooke’s law,Brekhovskikh1994 both the fiberglass coatingbudiansky1979pressure and the deployment conditions of the cable, e.g., burial depth and seafloor type,sladen2019distributed affect the strain sensitivity of the fiber-optic cable to the applied stress.
@@
DAS has gained momentum in geophysics as a cost-effective, high-resolution method for seismic monitoring and seismic tomography.mateeva2014distributed ; zhan2020distributed ; Lindsey2021 Applications range from near-surface seismic monitoring of ambient noisedou2017distributed ; saengduean2024multi to borehole measurements of hydraulic fracturingjin2017hydraulic ; chien2023automatic and monitoring the sea state dynamics with submarine cables.lindsey2019illuminating ; lin2024measurement ; chien2025calibration Recently, the potential of DAS technology has been leveraged for monitoring the underwater acoustic wavefield with promising results in the identification and localization of low-frequency sound sources such as blue and fin whale vocalizationsbouffaut2022eavesdropping ; landro2022sensing ; wilcock2023distributed ; Goestchel2025fin and ship noise.rivet2021preliminary ; shao2025tracking ; paap2025leveraging DAS converts the length of the fiber cable into an extensive linear aperture, which is especially useful for localization of low-frequency sound sources (below 100 Hz).rivet2021preliminary ; bouffaut2022eavesdropping ; landro2022sensing ; wilcock2023distributed Monitoring the ocean soundscape at low frequencieswilcock2014sounds is key to understanding the impact of anthropogenic noiseandrew2011long ; chapman2011low ; mcdonald2006increases and climate changegavrilov2008long on ecology,gavrilov2012steady security, and marine exploration.miksis2016low So far, the characterization of the ocean soundscape has been based on recordings from hydrophones and hydrophone arrays with limited spatial coverage.mcdonald2006increases ; chapman2011low
@@
Despite the increasing interest in DAS technology for high-resolution observations of acoustic signals, the current applications in ocean acoustics are limited to the aforementioned low-frequency wavefields, i.e., whale calls, ship noise, and ocean swell. The absence of ocean acoustic observations at higher frequencies is related to the choice of acquisition parameters of the current DAS systems, such as the interrogated fiber length and the pulse and gauge length, which are set with the objective of sufficient coverage and signal-to-noise ratio, respectively. Nevertheless, calibration experiments with controlled transmitted signalsdouglass2023distributed ; shen2024high ; rychen2023test and at-sea experiments for underwater communicationspotter2024first demonstrate that DAS detection capability is not limited to low-frequency sounds. This paper links the acquisition parameters with signal processing and provides insight for improved observations of acoustic signals.
@@
After trend removal and band-pass filtering to the frequency range of interest, the data are commonly analyzed with interferometric methods,dou2017distributed ; williams2022surface ; saengduean2024multi $f$-$k$ decomposition, and delay-and-sum beamforming.nasholm2022array Given the high dimensionality of DAS data, machine learning provides compelling methods for dimensionality reduction and feature clustering.chien2023automatic
@@
In this paper, we provide an overview of DAS technology. We detail the mathematical derivation of the underlying physics, from electromagnetic to mechanical and eventually acoustic quantities in Section II. The mathematical analysis highlights the physical principles of DAS in a generic manner, aiming to inform the choice of data acquisition parameters rather than to detail the technological specificity of different systems, which are evolving rapidly. Section III compares the spatial sensitivity of DAS to that of an array of spatially distributed sensors of a sizeable aperture. Section IV analyzes publicly available data for sound source detection based on the phase of the acoustic particle displacement. Finally, Section VII concludes the paper.
@@
$z$ axial distance $t$ time $k_{z}$ acoustic wavenumber along the fibre $\widetilde{[\cdot]}$ complex quantity
$L_{h}$ channel distance - distance between spatial samples $\tau$ two-way travel time $\theta$ incident angle $\cdot$$^{*}$ complex conjugation
$L_{p}$ light pulse length $T_{p}$ light pulse duration (width) $\mathbf{e}$ directional unit vector $\widehat{[\cdot]}$ estimated quantity
$L_{g}$ gauge length $\Delta\Phi$ differential phase of backscattered light $\cdot$$\cdot]$ convolution
$L_{f}$ interrogated fiber length $\tau_{\text{max}}$ two-way travel time along fiber length $\cdot]_{z}$ projection to the fiber’s axis
$f_{s}$ sampling frequency (spatial sampling) $f_{r}$ pulse repetition frequency (temporal sampling) $\cdot]_{p}$ quantity of the transmitted light pulse
$\cdot]_{b}$ quantity of the total backscattered light wave
$\cdot]_{a}$ quantity of the acoustic field
Figure 1: (color online) Schematic representation of scattering mechanisms due to a light pulse of wavelength $\lambda_{p}$ and frequency $f_{p}$ propagating through an optical fiber. Single-mode optical fibers are typically driven with laser sources at wavelengths $\lambda_{p}$ of about 1550 nm ($f_{p}=c/\lambda_{p}\approx$ 200 THz). External stimuli, such as an increase in temperature $T$ or strain $\epsilon$, affect the respective scattering response as indicated by the arrows. The horizontal axis represents the wavelength $\lambda$ or, in reverse direction, the frequency $f$ of the backscattered light. Stokes processes refer to an up-shift of the wavelength of the scattered light $\lambda>\lambda_{p}$, whereas anti-Stokes processes refer to a down-shift $\lambda<\lambda_{p}$. The Stokes radiation for Raman scattering is relatively insensitive to temperature and has a lower intensity than the corresponding anti-Stokes radiation due to lower scattering efficiency at lower frequencies.Ukil2011
@@ II Distributed Acoustic Sensing Technology
DAS measures the effect of external mechanical vibrations on the backscattering of laser light traveling along an optical fiber. This section details the DAS technology by providing an overview of the light scattering mechanisms within a fiber-optic cable (Sec. II.1), deriving the mathematical expression for the backscattered wave under monochromatic excitation (Secs. II.2-II.3), explaining the detection method of the backscattered light (Sec. II.4), showing how the measurand is related to strain from external mechanical stress (Sec. II.5), and finally explaining the typical acquisition parameters and the data structure (Sec. II.6). Table 1 provides a list of the important parameters involved.
@@ II.1 Scattering mechanisms
An optical fiber is a thin thread of silica glass with a core of dielectric material and cladding with a lower refractive index. The difference in refractive index between the core and the cladding creates an optical waveguide for light transmitted in the fiber. Optical fibers are designed to have a core with a constant refractive index such that light transmitted into the fiber travels undisturbed along the fiber. However, in practice, small density variations in the core material due to impurities or the manufacturing procedure introduce small random variations in the refractive index along the optical fiber, which cause scattering of the propagating electromagnetic wave.KennettBook Hence, the refractive index along the fiber’s axis can be described as
$n(z)=n_{f}+\mathrm{d}n(z),$ (1)
where $n_{f}$ is the nominal refractive index of the optical fiber’s core material and $\mathrm{d}n(z)$ describes the localized fluctuation of the refractive index due to material inhomogeneities. The refractive index determines the speed of light in the optical fiber as $c_{n}\approx c/n_{f}$, where $c\approx 3\times 10^{8}$ m/s is the speed of light in vacuum.
The main scattering mechanisms are Rayleigh, Brillouin, and Raman scattering.masoudi2016contributed Rayleigh scattering is an elastic process caused by local scatterers with dimensions much smaller than the wavelength that does not involve energy transfer to the scatterer and does not shift the frequency of the incident wave. Brillouin and Raman scattering are inelastic processes caused by changes in the vibrational states in a lattice and molecular level, respectively, that involve energy transfer to the scatterer and, as a result, shift the frequency of the incident wave.Cotter1983 ; Shiota1992 Brillouin scattering can be leveraged for both distributed strain and temperature measurements,Bao2011 whereas Raman scattering can be leveraged for distributed temperature sensing.Ukil2011 Rayleigh scattering exhibits the highest intensity and is the main scattering mechanism exploited for DAS measurements.KennettBook Figure 1 depicts the effect of external stimuli such as temperature and strain to elastic and inelastic optical scattering mechanisms [Grattan2000, , Ch. 3.8].
[…]
@@ II.2 Rayleigh backscattered wave
The active sensor in DAS technology is an optoelectronic system, referred to as the interrogation unit or, simply, the interrogator. The interrogator transmits a light pulse into the optical fiber and detects the intensity of the backscattered electric field due to the interaction of the transmitted pulse with the impurities along the fiber.Hartog2017
The most common single-mode optical fibers are designed with a cross-section of micro-metric dimensions such that only a single mode (transverse mode) of electromagnetic wave propagation is supported without modal dispersion. For a typical single-mode optical fiber, a minimum attenuation in the order of $0.2$ dB/km is achieved for a light wavelength of $1550$ nm [Hartog2017, , Fig. 2.3], i.e., at a frequency of $200$ THz. Hence, a narrowband laser pulse at this frequency would sustain a minimum transmission loss.Miah2017 ; Hartog2017
Multimode optical fibers have a core with a larger cross-section (at least an order of magnitude larger than single-mode fibers), which guides multiple electromagnetic modes simultaneously [Hartog2017, , Sec. 2.1]. Modal dispersion of the transmitted light as it propagates along the fiber eliminates its phase coherence and polarization information. Consequently, multimode optical fibers do not support highly sensitive detection methods, such as coherent optical reflectometry (see Appendix), resulting in reduced detection range and resolution. However, light emitting diodes (LED) are used for transmission in multimode fibers instead of highly coherent laser sources required in single-mode fibers, resulting in cost-effective systems [Grattan2000, , Ch. 1]. Despite the increased cost, single-mode fibers and coherent detection systems are the standard configuration for DAS due to the higher sensitivity.
Figure 2 depicts the mechanism of active optic sensing in single-mode fibers and indicates its utilization for passive sensing of external vibrations. The spatial dimension $z$ corresponds the axial distance from the interrogator, whereas the temporal dimensions $t=z/c_{n}$ and $\tau=2z/c_{n}$ correspond to one-way (unidirectional) and two-way (bidirectional) travel time of a light wave along the fiber, respectively.
In mathematical terms, a probe pulse [Fig. 2(b)] of duration $T_{p}$ and phase $\phi(t)$, generated by the laser source at $z=0$ is expressed as
$E_{p}(t)=p\left(\frac{t}{T_{p}}\right)\cos\left(\phi(t)\right).$ (2)
The pulse envelope $p$ is defined by the window function $w$ asharris2005use
@@
For example, for a rectangular window, $w(t)=1$. For mathematical simplicity, the window is considered symmetric around the origin. The pulse instantaneous phase is determined by the instantaneous angular frequency $\omega(t)$ as,
$\phi(t)=\int_{-T_{p}/2}^{t}\omega(\mathfrak{t})\mathrm{d}\mathfrak{t}.$ (4)
For a monochromatic continuous wave (CW), the pulse has a constant angular frequency $\omega(t)=\omega_{p}$, such that, with an initial phase $\phi_{0}$,
$\phi_{\text{CW}}(t)=\omega_{p}t+\phi_{0}.$ (5)
For a linearly frequency modulated chirped pulse (LFM), the angular frequency varies linearly with time within the pulse duration from $\omega_{1}$ to $\omega_{2}$ with central frequency $\omega_{p}=(\omega_{1}+\omega_{2})/2$ and bandwidth $BW=\omega_{2}-\omega_{1}$ such that $\omega(t)=\omega_{p}+\left(BW/T_{p}\right)t$ and the pulse phase is,
$\phi_{\text{LFM}}(t)=\left(\omega_{p}+\frac{BW}{2T_{p}}t\right)t+\phi_{0}.$ (6)
Figure 3 illustrates the temporal, spectral, and autocorrelation characteristics of continuous wave and frequency modulated pulses. Both continuous wave signalswilcock2023distributed and frequency modulated chirpswaagaard2021realtime are used as probe pulses with coherent detection methods in DAS technology. In particular, the inherent frequency diversity in frequency modulated probe pulses results in a correlation length defined by the bandwidth rather than by the pulse length for $BW\cdot T_{p}>1$. Hence, broadband pulses offer robustness against optical noise but require electronics (photodetectors/digitizers) with higher bandwidth increasing the complexity and cost of the system.Ruiz2019distributed
In the following, we adopt the monochromatic continuous wave model in Eq. (5) to simplify the mathematical derivations and to comply with the system used for the data analysis in later Sections.
Figure 3: Transmitted laser pulse characteristics. (a) Monochromatic pulse at angular frequency $\omega_{p}$ defined by a rectangular window of duration $T_{p}$, (b) its frequency spectrum, and (c) autocorrelation function. (d) Linear frequency modulated pulse of duration $T_{p}$ and bandwidth BWBW, (e) its frequency spectrum, and (f) autocorrelation function.
Consider the impulse response received at the interrogator at $z=0$ as a function of the arrival time $\tau$ due to an ideal point scatterer at $z_{s}$ as the two-way Green’s function
@@
where $\delta$ denotes the Dirac delta function, * is the convolution operator and $g_{i}$ and $g_{b}$ denote the Green’s function for the incident wave at location $z_{s}$ and time $t$ due to an impulse source at $z_{0}=0$ and $t_{0}=0$ (forward propagation) and the backscattered wave received at $z_{0}=0$ at time $\tau=2t$ due to an ideal scatterer at $z_{s}$ illuminated at time $t$ (backward propagation),
$g_{i}(t,z_{s}|t_{0},z_{0})\!=\!\delta\!\left(t-t_{0}-\frac{z_{s}\!-z_{0}}{c_{n}}\!\right)\!=\!\delta\!\left(t-\frac{z_{s}}{c_{n}}\right),$ (8)
$g_{b}(\tau,z_{0}|t,z_{s})\!=\!\delta\!\left(\tau-t+\frac{z_{0}\!-z_{s}}{c_{n}}\!\right)\!=\!\delta\!\left(t-\frac{z_{s}}{c_{n}}\right).$
More generally, the backscattered wave [Fig. 2(d)] from a point scatterer at $z_{s}$ with scattering strength $s(z_{s})$ due to the excitation pulse in Eq. (2) is
@@
External vibrations, such as an impinging acoustic wavefield at an incident angle $\theta$ characterized by particle displacement $\mathbf{u}$, apply stress to the fiber, which deforms its shape [Fig. 2(e)]. Under strain, any scatterer’s position is slightly perturbed by $\Delta z$, resulting in a proportional delay at the associated backscattered signal, while the scattering strength $s(z_{s})$ remains the same [Fig. 2(f)]: